OpenStudy (anonymous):

How does one solve the equation dy/dx - 4xy = 0 using an integrating factor? I see how to do it using a separation of variables approach, but how can an IF method work when the equation is set to 0?

sam (.sam.):

\[\frac{dy}{dx}-4xy=0\]

sam (.sam.):

Integrating factor is \[\mu(x)=e^{\int\limits P(x)dx}\] Where \[\frac{dy}{dx}+P(x)y=Q(x)\]

OpenStudy (anonymous):

But Q(x) is zero, so you end up with y=0

sam (.sam.):

Actually you don't need to, this is separable

OpenStudy (anonymous):

I know, but I have to do it using an integrating factor.

sam (.sam.):

Just do it, \[\mu(x)=e^{\int\limits-4xdx}=e^{-2x^2}\] Multiply across, \[e^{-2x^2}\frac{dy}{dx}-4e^{-2x^2}xy=0\] Then reverse product \[\frac{d}{dx}(e^{-2x^2}y)=0 \] Integrate both sides

sam (.sam.):

\[e^{-2x^2}y=c_1 \] \[y=c_1e^{2x^2}\]

OpenStudy (anonymous):

ah ok, that makes sense. cheers dude.

sam (.sam.):

yw

OpenStudy (anonymous):

@ParoxysmX ...proceed d same way and you will get it with a little bit rearrangement... or tell me how you did it and i will try to help..